OpenStudy (anonymous):

1/2(x-1/3)=-8/6x+6/6 need to solve symbolically

OpenStudy (anonymous):

\[\frac{1}{2} \left(-\frac{1}{3}+x\right)\text{ = }1-\frac{4 x}{3}\]Multiply each side by 6 and simplify.\[-1+3 x\text{ = }6-8 x\]Gather the x terms on the LHS and the other terms on the RHS and simplify.\[11x=7 \]

OpenStudy (anonymous):

how do i solve that symbolically < or>?